OpenStudy (anonymous):

HELP, ALG 2 DIVERGENT AND COVERGENT SERIES Identify whether the series summation of 12 open parentheses 3 over 5 close parentheses to the I minus 1 power from 1 to infinity is a convergent or divergent geometric series and find the sum, if possible. This is a convergent geometric series. The sum cannot be found. This is a convergent geometric series. The sum is 30. This is a divergent geometric series. The sum cannot be found. This is a divergent geometric series. The sum is 30.

OpenStudy (anonymous):

@ShadowLegendX PLEASE HELP

OpenStudy (anonymous):

Convergent series have sums, and divergent series don't have sums, so you can eliminate the first and last options.

OpenStudy (anonymous):

Is the series \[\large\sum_{i=1}^\infty12\left(\frac{3}{5}\right)^{i-1}~~?\]

OpenStudy (anonymous):

If it is... recall that a geometric series with ratio \(r\) converges for \(|r|<1\), or \(-1<r<1\). Does \(r=\dfrac{3}{5}\) fall between -1 and 1? If yes, the series converges. If no, it diverges.

OpenStudy (anonymous):

If you think it converges, you can use this formula to find the sum: \[\large\sum_{i=1}^\infty ar^{i-1}=\frac{a}{1-r}\]